# Aligning ASR Evaluation with Human and LLM Judgments: Intelligibility Metrics Using Phonetic, Semantic, and NLI Approaches

### v1

@@ 1 Introduction
Two key challenges emerge from these observations. First, existing ASR metrics do not sufficiently capture intelligibility. Alternative metrics such as BERTScore [1] generally offer semantic similarity but may fail to detect logical contradictions in certain cases. For example, it assigns a high similarity score (0.8741) between the following contradictory sentences:
Hypothesis: ”The project deadline is flexible” vs. Reference: ”The project deadline is not flexible”.
@@
Second, while LLMs can improve ASR outputs, their correction capability remains underexplored. Recent advancements suggest that LLMs can refine ASR-generated transcripts[2], [3], [4], [5], correcting phonetic errors and improving intelligibility. However, it is unclear which types of errors LLMs can correct most effectively and how this aligns with human perception. Understanding the correctability of ASR transcripts can inform better ASR evaluation methods.
@@
Using data from the Speech Accessibility Project (SAP), we analyze multiple metrics while assessing several ASR systems across varying severity levels of dysarthric speech. Leveraging insights from this analysis, our optimized metric achieves a high correlation (0.890) with human judgments, notably outperforming traditional metrics.
@@ 2 Challenges in ASR evaluation for dysarthric speech
We evaluate multiple ASR systems, including Wav2vec 2.0 [6] and Whisper [7], with dysarthric speech from the validation split (23.32h) of SAP [8] release 2023-10-05 for this investigation. The experiments are conducted solely on SAP, as it is currently the largest and most diverse publicly available corpus for dysarthric speech, approximately ten times larger than any other comparable dataset, making it the most reliable benchmark for evaluating ASR performance on impaired speech. Wav2vec 2.0 base is a transformer-based speech encoder pre-trained with Librispeech, which performs well in ASR by fine-tuning on a specific corpus to optimize the CTC [9] loss. This work utilized two official wav2vec base models, fine-tuned on typical speech from $100$ hours and $960$ hours of Librispeech, respectively, as well as the best checkpoint trained by [10], using $151.47$ hours of dysarthric speech in the SAP-1005 training split. We refer to these models as wav2vec-(libri100, libri960 and sap1005) throughout the paper. Whisper base English model utilized in this work is a multi-task multi-lingual transformer-based encoder-decoder model trained with 680,000 hours of audio using weak supervision, which significantly enhances its accuracy and robustness in ASR across different scenarios.
We evaluate the ASR systems using three metric categories: word-level, phonetic, and semantic similarity. For word-level similarity, we use WER, calculating it for each utterance and averaging the results to maintain consistency. Phonetic similarity is assessed through g2p [^1] https://github.com/Kyubyong/g2p phoneme-level similarity and Soundex code-based similarity [^2] https://ics.uci.edu/~dan/genealogy/Miller/javascrp/soundex.htm. For semantic similarity, we apply BERTScore F1 [1] and BLEURT [11]. Additionally, we use Heval [12], a hybrid metric developed specifically for ASR systems, to examine how these metrics perform with dysarthric speech.
@@
Table 1: Evaluation of ASR Systems on Dysarthric Speech using WER, Phonetic, and Semantic Metrics. “Psim” refers to “phonetic similarity”, calculated by Levenshtein Distance (Psim I) and Soundex Algorithm (Psim II). The best entry in each column is in boldface.
Model WERPsim I Psim II BERT Bleurt Heval
@@
Table 2 shows the ASR performance across four severity levels of dysarthric speech: high (2h), medium (4h), low (7h), and very low (8h), based on [10]. In this table, we consider only Soundex (Psim II) for phonetic similarity, as both g2p and Soundex provide nearly the same interpretation, with Soundex offering better insight into phonetic variations. Wav2vec-sap1005 outperforms Whisper at high and medium severity levels, with WER differences of 16.89% (51.50% vs. 68.39%) and 9.45% (29.84% vs. 39.29%). However, Whisper performs better at low and very low severity levels, with advantages of 1.61% (19.64% vs. 21.25%) and 4.02% (10.72% vs. 14.74%). This suggests Whisper generalizes better to low-severity speech, while wav2vec-sap1005 excels in severe cases. Consistent with Table 1, Whisper achieves the highest semantic similarity across most severity levels, except for high severity.
@@
Model WERPsim II BERT Bleurt Heval
@@ H Wav2vec
Table 3: Evaluation of ASR Systems on Dysarthric Speech, w/ and w/o GPT corrections. “Wav2vec” in this table refers to wav2vec-sap1005. The best entry in each block is in boldface.
Model WERPsim II BERT Bleurt Heval
@@ 3 Evaluating ASR Correctability Using Large Language Models
While WER, phonetic similarity, and semantic similarity capture different aspects of ASR quality, they do not directly indicate how intelligible an ASR output is or how easily it can be corrected. Much recent research has demonstrated that LLM correction can reduce the WER of an ASR transcript [2, 3, 4, 5, 13], but the aspects of a transcript that make it correctable or uncorrectable have not been fully explored. In this section, we evaluate whether phonetic similarity, semantic similarity, and WER contribute to predicting ASR correctability. To investigate this, we tested two LLM-based correction settings: (1) the WavLLM model [14], which integrates LLMs with Whisper [7] and WavLM [15] for ASR tasks, and (2) the GPT-4 API [16], using a prompt to generate corrected transcripts. GPT-4 is chosen for its reliability over LLaMA3 and GPT-3, and we limited the analysis to 1,441 samples from the SAP1005 dataset (1,321 short, 50 moderate, and 50 long samples), using a model temperature of 0.1. The transcript lengths are categorized by word-level tokens: short (fewer than 10), moderate (10 to 25), and long (more than 30).
Prompt: ”Correct this ASR transcript for readability, clarity, and spelling while preserving the original meaning. Make minimal changes, fixing errors or incorrect terms and names without unnecessary rephrasing.”
@@
  • Hypothesis: “SET THE AIR CONDITIONING CONDITI CONDITIONING TO SEESEVENTY SE E EE SEENT EIGEIGHT”
  • Reference: “SET THE AIR CONDITION DITIONING TO SEV SEVENTY EIGHT”
  • GPT-Corrected: “SET THE AIR CONDITIONING CONDITIONING TO SEVENTY EIGHT”
@@
  • Hypothesis: “OPEN GULAMNBA”
  • Reference: “OPEN DUOLINGO”
  • GPT-Corrected: “OPEN GULAMNBA CORRECTED TEXT OPEN GYM NBA”
In this case, GPT-4 completely misinterprets the intended phrase ”DUOLINGO” and introduces an unrelated correction, resulting in an even worse transcription. The WER increases significantly from 0.5 to 3.0, demonstrating a severe case of hallucination.
@@ 4 A new integrated metric for dysartic speech ASR evaluation
Calculating NLI_Score: An ASR hypothesis might be considered to correctly represent user intent if it can be proven to logically entails the reference transcript, even if the semantic and phonetic similarities between reference and hypothesis are quite low. Entailment probabilities computed by a trained NLI model have been used to estimate the quality of machine translation and document summarization [17] by computing a weighted linear average of the probability that the hypothesis entails the reference (in the notation of [17]: $P_{e}(\text{ref}\leftarrow\text{hypo})$), the probability that the hypothesis entails the reference ($P_{e}(\text{hypo}\leftarrow\text{ref})$), the probability of mutual entailment ($P_{e}(\text{hypo}\leftrightarrow\text{ref})$), the probabilities of all three types of contradiction ($P_{c}$), and the probability that there is neither entailment nor contradiction ($P_{n}$). Of these, the scalar probability $P_{e}(\text{hypo}\leftrightarrow\text{ref})$ was demonstrated to be the most effective metric of machine translation and document summarization quality [17], and we therefore adopt the same metric as our NLI score ($S_{NLI}$) of ASR hypothesis quality. $S_{NLI}$ is set equal to the probability $P_{e}(\text{hypo}\leftrightarrow\text{ref})$ computed by a RoBERTa-large model [18] that has beeen fine-tuned on the SNLI [19], MNLI, FEVER [20], and ANLI [21] datasets. To calculate the semantic score $S_{S}$, we used BERTScore, and for phonetic similarity $S_{P}$, we applied Soundex with Jaro-Winkler similarity, as discussed in Section 2. The integrated metric is defined as:
[…]

### v2

@@ 1 Introduction
Two key challenges emerge from these observations. First, existing ASR metrics do not sufficiently capture intelligibility. Alternative metrics such as BERTScore [zhang2020bert] generally offer semantic similarity but may fail to detect logical contradictions in certain cases. For example, it assigns a high similarity score (0.8741) between the following contradictory sentences:
Hypothesis: "The project deadline is flexible" vs. Reference: "The project deadline is not flexible".
@@
Second, while LLMs can improve ASR outputs, their correction capability remains underexplored. Recent advancements suggest that LLMs can refine ASR-generated transcripts[guo2019spelling], [shivakumar2019learning], [tanaka2018neural], [gu2024denoising], correcting phonetic errors and improving intelligibility. However, it is unclear which types of errors LLMs can correct most effectively and how this aligns with human perception. Understanding the correctability of ASR transcripts can inform better ASR evaluation methods.
@@
Using data from the Speech Accessibility Project (SAP), we analyze multiple metrics while assessing several ASR systems across varying severity levels of dysarthric speech. Leveraging insights from this analysis, our optimized metric achieves a high correlation (0.890) with human judgments, notably outperforming traditional metrics. The evaluation scripts are available at the GitHub repository [^1] https://github.com/BornaliP/SemScore.
@@ 2 Challenges in ASR evaluation for dysarthric speech
We evaluate multiple ASR systems, including Wav2vec 2.0 [baevski2020wav2vec] and Whisper [radford2022robust], with dysarthric speech from the validation split (23.32h) of SAP [omalley2023uiuc] release 2023-10-05 for this investigation. The experiments are conducted solely on SAP, as it is currently the largest and most diverse publicly available corpus for dysarthric speech, approximately ten times larger than any other comparable dataset, making it the most reliable benchmark for evaluating ASR performance on impaired speech. Wav2vec 2.0 base is a transformer-based speech encoder pre-trained with Librispeech, which performs well in ASR by fine-tuning on a specific corpus to optimize the CTC [graves2006connectionist] loss. This work utilized two official wav2vec base models, fine-tuned on typical speech from $100$ hours and $960$ hours of Librispeech, respectively, as well as the best checkpoint trained by [Zheng2024], using $151.47$ hours of dysarthric speech in the SAP-1005 training split. We refer to these models as wav2vec-(libri100, libri960 and sap1005) throughout the paper. Whisper base English model utilized in this work is a multi-task multi-lingual transformer-based encoder-decoder model trained with 680,000 hours of audio using weak supervision, which significantly enhances its accuracy and robustness in ASR across different scenarios.
We evaluate the ASR systems using three metric categories: word-level, phonetic, and semantic similarity. For word-level similarity, we use WER, calculating it for each utterance and averaging the results to maintain consistency. Phonetic similarity is assessed through g2p [^2] https://github.com/Kyubyong/g2p phoneme-level similarity and Soundex code-based similarity [^3] https://ics.uci.edu/~dan/genealogy/Miller/javascrp/soundex.htm. For semantic similarity, we apply BERTScore F1 [zhang2020bert] and BLEURT [sellam2020bleurt]. Additionally, we use Heval [sasindran2023h], a hybrid metric developed specifically for ASR systems, to examine how these metrics perform with dysarthric speech.
@@
Table 1: Evaluation of ASR Systems on Dysarthric Speech using WER, Phonetic, and Semantic Metrics. ``Psim'' refers to ``phonetic similarity'', calculated by Levenshtein Distance (Psim I) and Soundex Algorithm (Psim II). The best entry in each column is in boldface.
Model WER% Psim I Psim II BERT Bleurt Heval
@@
Table 2 shows the ASR performance across four severity levels of dysarthric speech: high (2h), medium (4h), low (7h), and very low (8h), based on [Zheng2024]. In this table, we consider only Soundex (Psim II) for phonetic similarity, as both g2p and Soundex provide nearly the same interpretation, with Soundex offering better insight into phonetic variations. Wav2vec-sap1005 outperforms Whisper at high and medium severity levels, with WER differences of 16.89% (51.50% vs. 68.39%) and 9.45% (29.84% vs. 39.29%). However, Whisper performs better at low and very low severity levels, with advantages of 1.61% (19.64% vs. 21.25%) and 4.02% (10.72% vs. 14.74%). This suggests Whisper generalizes better to low-severity speech, while wav2vec-sap1005 excels in severe cases. Consistent with Table 1, Whisper achieves the highest semantic similarity across most severity levels, except for high severity.
@@
Model WER% Psim II BERT Bleurt Heval
@@ H Wav2vec
Table 3: Evaluation of ASR Systems on Dysarthric Speech, w/ and w/o GPT corrections. ``Wav2vec'' in this table refers to wav2vec-sap1005. The best entry in each block is in boldface.
Model WER% Psim II BERT Bleurt Heval
@@ 3 Evaluating ASR Correctability Using Large Language Models
While WER, phonetic similarity, and semantic similarity capture different aspects of ASR quality, they do not directly indicate how intelligible an ASR output is or how easily it can be corrected. Much recent research has demonstrated that LLM correction can reduce the WER of an ASR transcript [guo2019spelling, shivakumar2019learning, tanaka2018neural, gu2024denoising, tomanek2024large], but the aspects of a transcript that make it correctable or uncorrectable have not been fully explored. In this section, we evaluate whether phonetic similarity, semantic similarity, and WER contribute to predicting ASR correctability. To investigate this, we tested two LLM-based correction settings: (1) the WavLLM model [hu2024wavllm], which integrates LLMs with Whisper [radford2022robust] and WavLM [chen2022wavlm] for ASR tasks, and (2) the GPT-4 API [achiam2023gpt], using a prompt to generate corrected transcripts. GPT-4 is chosen for its reliability over LLaMA3 and GPT-3, and we limited the analysis to 1,441 samples from the SAP1005 dataset (1,321 short, 50 moderate, and 50 long samples), using a model temperature of 0.1. The transcript lengths are categorized by word-level tokens: short (fewer than 10), moderate (10 to 25), and long (more than 30).
Prompt: "Correct this ASR transcript for readability, clarity, and spelling while preserving the original meaning. Make minimal changes, fixing errors or incorrect terms and names without unnecessary rephrasing."
@@
  • Hypothesis: ``SET THE AIR CONDITIONING CONDITI CONDITIONING TO SEESEVENTY SE E EE SEENT EIGEIGHT''
  • Reference: ``SET THE AIR CONDITION DITIONING TO SEV SEVENTY EIGHT''
  • GPT-Corrected: ``SET THE AIR CONDITIONING CONDITIONING TO SEVENTY EIGHT''
@@
  • Hypothesis: ``OPEN GULAMNBA''
  • Reference: ``OPEN DUOLINGO''
  • GPT-Corrected: ``OPEN GULAMNBA CORRECTED TEXT OPEN GYM NBA''
In this case, GPT-4 completely misinterprets the intended phrase "DUOLINGO" and introduces an unrelated correction, resulting in an even worse transcription. The WER increases significantly from 0.5 to 3.0, demonstrating a severe case of hallucination.
@@ 4 A new integrated metric for dysartic speech ASR evaluation
Calculating NLI_Score: An ASR hypothesis might be considered to correctly represent user intent if it can be proven to logically entails the reference transcript, even if the semantic and phonetic similarities between reference and hypothesis are quite low. Entailment probabilities computed by a trained NLI model have been used to estimate the quality of machine translation and document summarization [chen2023menli] by computing a weighted linear average of the probability that the hypothesis entails the reference (in the notation of [chen2023menli]: $P_{e}(\text{ref}\leftarrow\text{hypo})$), the probability that the hypothesis entails the reference ($P_{e}(\text{hypo}\leftarrow\text{ref})$), the probability of mutual entailment ($P_{e}(\text{hypo}\leftrightarrow\text{ref})$), the probabilities of all three types of contradiction ($P_{c}$), and the probability that there is neither entailment nor contradiction ($P_{n}$). Of these, the scalar probability $P_{e}(\text{hypo}\leftrightarrow\text{ref})$ was demonstrated to be the most effective metric of machine translation and document summarization quality [chen2023menli], and we therefore adopt the same metric as our NLI score ($S_{NLI}$) of ASR hypothesis quality. $S_{NLI}$ is set equal to the probability $P_{e}(\text{hypo}\leftrightarrow\text{ref})$ computed by a RoBERTa-large model [liu2019roberta] that has beeen fine-tuned on the SNLI [bowman-etal-2015-large], MNLI, FEVER [nie2019combining], and ANLI [nie2019adversarial] datasets. To calculate the semantic score $S_{S}$, we used BERTScore, and for phonetic similarity $S_{P}$, we applied Soundex with Jaro-Winkler similarity, as discussed in Section 2. The integrated metric is defined as:
$\text{Integrated\_Score}=\alpha S_{NLI}+\beta S_{S}+\gamma S_{P}$